OpenStudy (anonymous):

i need someone to show me how to work the problem (s-3)(3s+6)=0

OpenStudy (anonymous):

You can split this equation into two, because we want to know where 0=0 s - 3 = 0, so s = 3 3s + 6 = 0 s = -2

OpenStudy (anonymous):

When S = 0, equate the equation. So (s-3) or (3s+6) = 0, because a number multiplying zero will be zero. So S = 3 or 3S = 6 for either of the bracketed numbers to be zero.

OpenStudy (anonymous):

ok ty both very much that helps and makes total sense my brain wasnt comprehending lol

OpenStudy (anonymous):

It's alright. Practice more and you'll be able to do it mentally. All the best. :)